OpenStudy (anonymous):

]Choose the equation of the line passing through the point (2, -4) and parallel to y = 3x - 6. y = 3x - 10 y = 3x - 6 y = 3x - 2 y = 3x - 12

OpenStudy (anonymous):

Same slope of 3 y = 3x + b Plug in known point (2, −4) -4 = 3(2) + b -4 = 6 + b b = −10 Now have m= 3 and b = 2 y = 3x - 10

OpenStudy (anonymous):

Do you understand?

OpenStudy (anonymous):

Thanks ;). Yes thank yo

OpenStudy (anonymous):

You welcome